OpenStudy (anonymous):

what is 1 foot and 3.5 inches as a fraction

OpenStudy (chrisdbest):

If 1 is represented as a whole number, then you baseline is 1

OpenStudy (mathstudent55):

1 ft = 12 in.

OpenStudy (chrisdbest):

3.5 inches= .29 ft

OpenStudy (chrisdbest):

So 1ft 3.5"= 1.29 ft

OpenStudy (anonymous):

what they said ^

OpenStudy (anonymous):

thx guys

OpenStudy (mathstudent55):

1 ft and 3.5 in. = 12 in. + 3.5 in. = 15.5 in. \(15.5 ~in. \times \dfrac{1~ft}{12~in.} = \dfrac{15.5}{12} ~ft\) Now you can reduce the fraction.

OpenStudy (chrisdbest):

so \[\frac{ 129 }{ 100 }\] is a fraction form

OpenStudy (mathstudent55):

\(\dfrac{129}{100}\) is a fraction, but it is not the correct answer to this problem. It's only an approximation.

OpenStudy (mathstudent55):

\(\dfrac{15.5}{12} = \dfrac{155}{120} = \dfrac{31}{24} \) This is an exact answer.